OpenStudy (anonymous):

Which factor do you cancel from the numerator and denominator to simplify ? x^2 + 4x - 5 / x^2 - 25

OpenStudy (sburchette):

If you factor the numerator and denominator you will get\[\frac{(x-1)(x+5)}{(x+5)(x-5)}\] Because there is an (x+5) in both the numerator and denominator, they will cancel each other and you will be left with \[\frac{x-1}{x-5}\]

OpenStudy (mimi_x3):

\[\large \frac{x^{2}+4x-5}{x^{2}-25} \] First one: \(x^2+4x-5\) Trinomials: \(x^2+x(a+b)+ab\) Second one: \(x^2-25\) Difference of two squares: \((a+b)(a-b)=a^2-b^2\)

OpenStudy (anonymous):

Multiple choice lol x + 4 x - 4 x - 5 x + 5

OpenStudy (mimi_x3):

Well, which one do you think is right?

OpenStudy (mimi_x3):

wait..the choices are wrong..

OpenStudy (anonymous):

That's what it looked like to me but thats exactly what they are